Main INDEX
Monthly INDEX
PREV
NEXT
User name cusanno
Log entry time 10:02:01 on August 05, 2006
Entry number 177565
Followups:
keyword=target cell pressure increases
In the last two hour pressure of both target cells are increasing.
Pressure went from 28 to 29 on LH2 and from 29 to 30 on LD2. Present beam
is 4 microA on LH2.
I do not thing this is critical, just an observation. I will check what
happens during the day.
A copy of this log entry has been emailed to: meekins@jlab.org